OpenStudy (anonymous):

determine whether the sequence [an] diverges or converges an = squareroot of (n^2 − 2n + 8) -n i've tried simplifying by multiplying numerator and denominator by its conjugate but i still got divergent. it converges to some number but i don't know what − n

OpenStudy (amistre64):

for larger and larger values of n, this begins to resemble sqrt(n^2)-n the question i would have is how fast is it doing that?

OpenStudy (amistre64):

are you using a denominator of 1? its hard to see if im reading your post correctly

OpenStudy (anonymous):

yes it's a denominator of 1.

OpenStudy (anonymous):

it's just the (squareroot of n^2 -2n +8) -n

OpenStudy (tkhunny):

Try this: \(\dfrac{\sqrt{n^{2}-2n+8}-n}{1}\cdot\dfrac{\sqrt{n^{2}-2n+8}+n}{\sqrt{n^{2}-2n+8}+n}\). Simplify numerator and denominator separately. Something magical may occur.

OpenStudy (anonymous):

(-2n+8) / denominator. . I factored out a 'n' from numerator and from the squareroot. Then I reduced the fraction by canceling the n's. I took the limit of n to infinity and got #/0. :/

OpenStudy (tkhunny):

How did you manage 0 in the denominator? That's no good.

OpenStudy (anonymous):

Oh oops. I see my mistake. I subtracted n instead of adding it

OpenStudy (anonymous):

Thanks!

OpenStudy (tkhunny):

Did you get the non-zero limit?

OpenStudy (anonymous):

Yes -1

OpenStudy (tkhunny):

Perfect. Tricky, wasn't it. Did you use l'Hopital?

OpenStudy (anonymous):

No

OpenStudy (tkhunny):

That's good, since it would be a wonderous circular discouragement. It's a funny case for l'Hopital. Good work reasoning it out without that confusion.